OpenStudy (anonymous):

factor 18x3y-15x2y2-18xy

OpenStudy (anonymous):

common factor of \[3xy\] to start

OpenStudy (anonymous):

3xy(x^2-5xy-6)

OpenStudy (anonymous):

18x3y-15x2y2-18xy^3 =3xy(6x^2-5xy-6y^2) =3xy(6x^2-9xy+4xy-6y^2) =3xy(3x(2x-3y)+2y(2x-3y)) =3xy(3x+2y)(2x-3y)